Please help with budget (wbs)


Costs per 30 days  Transport costs per school is R50 000   Total costs of Materials (Meals, Mahewu and Contraceptives) per school R 275 000  Director’s salary is R 120 000 (spread over 5 schools)   Social worker’s salary is R25 000( spread over 5 schools)  Total costs of catering staff wages is R 15 000 per school  Total cost of catering equipment is R25 000 per school  Project office costs is 12% of the sub-total  Profit is 24% of the sub-total.  


1
Expert's answer
2021-09-03T04:52:02-0400

For 60 days:-

Transport costs per school =R50000*5*2= R 5 00 000

Total costs of Materials (Meals, Mahewu and Contraceptives) per school R 275 000=R 2 75 000*2=R 5 50 000

Director's salary is R 120 000 (spread over 5 schools)= R 1 20 000*5*2=R 12 00 000 

Social worker's salary is R25 000( spread over 5 schools)= R 25 000*3*2=R 1 50 000 

Total costs of catering staff wages is R 15 000 per school= R 15 000*5*2= R 1 50 000

Total cost of catering equipment is R25 000 per school= R 25 000*5*2= R 25 00 000

                                                                                                                  Total= R 50 50 000 

Project office costs is 12% of the sub-total= 12* R 50 50 000/100= R 6 06 000

Total cost incurred = R 50 50 000 + R 6 06 000=R 56 56 000

Profit is 24% of the sub-total= 24* R 56 56 000/100= R 13 57 440

 

In Latex code

Total cost= Transport costs per school =R50000\times 5\times 2= R 5 00 000

Total costs of Materials (Meals, Mahewu and Contraceptives) per school R 275 000=R 2 75 000\pm \times 2=R 5 50 000

Director's salary is R 120 000 (spread over 5 schools)= R 1 20 000\times 5\times 2=R 12 00 000 

Social worker's salary is R25 000( spread over 5 schools)= R 25 000\times 3\times 2=R 1 50 000 

Total costs of catering staff wages is R 15 000 per school= R 15 000\times 5\times 2= R 1 50 000

Total cost of catering equipment is R25 000 per school= R 25 000\times 5\times 2= R 25 00 000

                                                                                                                 Total= R 50 50 000 

 

Project office costs is 12% of the sub-total= 12\times  R 50 50 000\div 100= R 6 06 000

Total cost incurred = R 50 50 000 \pm  R 6 06 000=R 56 56 000

Profit is 24% of the sub-total= 24\times  R 56 56 000/100= R 13 57 440

 

 

Project director works for spreading the educationand social awareness with the respective departments for objective attainment. The main objective is to curb early pregnancy.